OpenStudy (anonymous):

If one root of the equation \(8x^2 - 4x + p = 0\) is six times the other, then \(p\) is equal to: A. 0 B. 15/49 C. 13/49 D. 12/49

OpenStudy (anonymous):

This question is from 10th grade A.P.

OpenStudy (anonymous):

Here's my approach: Let the roots be \(a\) and \(6a\). The sum of roots is \(1/2\). So \(7a = 1/2\), therefore \(a = 1/14\) and the other root is \(6/14\). The product of roots is \(p/8= 3/98\) then \(p=12/49\).

OpenStudy (anonymous):

Yeeeeee, solved it!